Commissioning Notes from 16-17 May 2008 MST

Observers: Doug M on mountain. John H and Andrew R in Tucson. T.O.: Randy G.

Summary

After a few stops and starts, collected the data we needed to correct rotator off-set, and coefficients for z5,6,7,8 and 11. Will analyze and try out results tomorrow night. Fault on EL axis seemingly down to some loose wires. Weather was clear all night with seeing hovering around 1.5".

Nighttime Activities

3:25 Setting up. Clouds but no rain at the moment.

3:40 Setting up. Need to run "prep.dir" in IRAF to save data.

3:41 Elevation axis not slewing to track. Error message in MCS "RR Power Amps not normal, power amplifier has falied". Fixed, restarted the elevation drive.

4:01 New offset for pointing. go to lbtmu102, "ptlist" lists elements in pointing model. modify "IE", elevation zero point, and "CA", azimuth optical offset. To change, for example "ptmodify CA #", # in arc-seconds.

4:20 Center star on IRTC with CA = -93 IE = -42.

4:33 Acquired star on guider

5:15 Achieved guiding lock, attempting to get wfs image, camera freezing, ctrl c not working, stopped and restarted AGw

5:27 sent point to command, got a success message but no guiding. Stopped and restarted GCSL

5:28 Found IRS had stopped, but IRAF routine reported success.

5:30 humidity 88%, mostly clear, seeing 1.5" according to Polaris monitor

5:41 Problem with WFS, spot images very large, seeing affected. WFS not returning sensible slopes. See wfsc000010_dd.fits in /Repository/AGW_Data/080517w

5:51 Obtained useful results with wfs, sending focus commands to psf with +1 coefficient

5:51 GCS unexpectedly dies, stop GCSL

5:57 Three attempts to restart guiding, third time the charm...

6:04 Sent focus correction to PSF, worked well with Z4 coeffcient +1 worked well, sent z6 with -1 coefficient, went the wrong way by the same amount, changed sign of coefficient, went to resend and lost guiding. El axis stopped tracking.

6:15 cleared forces "clear C00" to start again

6:24 using z4 coeff = +1, z6 coeff +1 got lots of z8, found a problem with turing off zernike terms, back to drawing board, removed active optics

6:36 El drive stopped for thrid time, Randy will produce issue track. Have decided to clear active optics, re-focus, and then obtain multiple data sets in the following way: set z4 to +1000, take 10 WFS images, set all terms to zero, take 10 images, set z5 to 1000nm, take ten images, set z5 to zero, take ten images, and so on up to z11. Will then post-analyze to determine coefficients and rotation offset.

6:55 guider trouble. Turns out OAC server not working properly. Restarted AGW.

7:07 AGW UMAC is having difficulties, loses communication and then reloads codes ~ every 4th readout, i.e. every ~ 20 seconds. See issuetrak 1297.

7:16 Achieved guiding lock, cleared active optics, have about 1000 nm z4.

7:21 set focus in psf window to -0.069 mm taking block of 10 WFS images 30 second exposures, 1.6" seeing, cleared active optics apart from focus. 10 images from WFSC000035-44.fits

7:35 focus drifted quite a lot, reset focus to -0.043 mm in PSF GUI

7:37 set z5 to +1000 take images from WFSC000046-55

7:38 created a file WFSC000035-44.txt, a file with zern coefficients from block of 10 observations mentioned above

7:42 lost guiding last good wfsc image WFSC000050 elevation drive carcks it for the 5th time, Randy goes upstairs to powercycle drives

7:44 created a file WFSC000046-50.txt, a file with zern coefficients from block of 10 observations mentioned above

7:51 cleared active optics, reset focus to -0.043 mm in PSF GUI

7:58 Randy returns, power cycling required two attempts, see Randy's issue trak

8:30 El drives fail again, attempting troubleshoot with Jon R.

9:25 Randy returns, back on sky, getting guiding set up.

9:45 acquired guide star, reset exposure time on guide cam to 200 ms

9:49 Trying to adjust focus to zero from ~ 500. Sending 300 nm in PSF gui and focus doesn't change, sent 500 nm and it moved 500! according to PSF gui the mirror was moving, according to WFS it didn't move until we told it to move 500 nm of Z4...

9:53 taking 5 wfs images with cleared active optics (apart from Z4) wfsc000056-60

9:54 GCS died unexpectedly restart GCSL, tracking was good enough unguided for our images

10:05 issue trak for elevation axis problem is 1360

10:15 lost guiding, restart GCSL

10:19 taking 5 images with 1000 nm z6 (wfsc000069-73) check coefficient against Z5.

10:24 taken 5 images with active optics cleared (wfsc000075-79)

10:36 taken 5 image with 1000 nm z7 (wfsc000083-87)

10:40 lost guiding , then got it again without doing anything

10:43 take five images with cleared active optics (wfsc000089-93)

10:48 cleared active optics, refocussed and -1000 nm z8 take 5 images (wfsc000094-98)

10:58 John H, help please, we hit an x limit and don't know how to put in a pointing rotation in PSF to get rid of it.

10:59 cleared active optics (with a little coma left on to keep us off the x limit) refocused, take 5 images (wfsc00099-103)

11:04 add 1000 nm z11, too aberrated for WFS. Also get very large z4 changes with z11 changes. Have recent changes from Chris B affected the amount of mirror piston going into the z11 term?

11:20 taking 5 images with 1600 nm of focus when we put in -500 nm of Z11 (wfsc000110-114)

11:28 taking 5 images with new focus adjustment to get rid of the 1600 nm focus above, z11 still set to -500 (wfsc000118-122)

11:33 reset global offset in x to 1.9 to stop hitting x limit

11:39 guider imager seems to be oscillating

11:40 taking five images at nominal zero (wfsc000126-130) washed out by brightening sky...

-- AndrewRakich - 17 May 2008
Topic revision: r3 - 18 May 2008, JohnHill
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback